# Signal Processing for Haptic Surface Modeling: a Review

### v1

@@ 6 Research tasks
Dataset Single mode Multi- modal Cross- modal Cross-modal generation Stylization Grasp prediction Contact localization Perceptual similarity Video generation/ prediction
HaTT [142] [143] [144] [145] [146, 147] [148]
LMT [109] [111][149] [150, 151, 152] [153, 154] [155, 156] [157] [158, 159] [160, 161] [162]
FoS [112][163] [164, 165] [166, 167]
Fabrics [168, 169] [149, 152]
Texplorer [120] [115, 120] [170, 171] [143, 172] [157] [173] [174]
ViTac [116] [175, 176] [176] [116][176] [177]
MTaF [117][178]
Haptex [179]
VisGel [180] [181] [180]
ObjectFolder [125][182] [121][125] [180, 163] [125] [125][163] [125] [180]
Touch and Go [178, 183] [163] [108][178] [183, 184] [163, 180] [181, 163] [162] [108] [163, 181] [108][163] [108][180]
PoseIt [123]
SSVTP [163] [124] [163] [163] [124]
MPI-10 [185]
SSVTP + HCT [129] [129]
TaRF [132] [132] [132]
@@
In this chapter we provide an overview of the main tasks currently tackled by researchers in the haptic signal processing domain. Starting from the datasets listed in Section 5, in Tab. 4 we provide a thorough classification of the available works for each task. While some of them, like classification and compression, are akin to other domains and straightforward to understand, other tasks are very specific to haptics, like contact localisation and grasp prediction. For a comprehensive overview, in the next paragraphs we delve into each task, highlighting their peculiarities as well as the common metrics used for the evaluation.
@@
is one of the tasks mainly performed when dealing with haptic data. As the collected datasets usually provide ground truths in different domains (e.g. haptic maps, RGB images, audio produced by the object upon contact) for the same sample, classification can be performed according to different strategies: single mode, multi-modal, and cross-modal. Works adopting the single mode strategy aim to classify materials/objects by training and testing in the same domain. For approaches employing only haptic data, this task is usually performed on tactile maps [116, 120, 168, 175, 176, 121, 172, 169, 125, 178, 183, 163], since dealing with images enables the usage of established approaches, such as deep neural networks (DNNs). Differently, other works, such as [109, 142] proposed a material classification method that exploits acceleration data, achieving state-of-the-art results and showing that acceleration data establish a unique fingerprint for each material. Multi-modal approaches employ multiple data sources to classify materials and objects. These works rely mainly on visual and haptic data [150, 176, 152, 151, 124, 149, 108, 153, 182, 143, 178, 132, 129, 184, 163], since their combination allows for the exploitation of both mono-dimensional and multi-dimensional data. However, also audio cues are widely used jointly with haptic data [154, 111, 115, 171, 120, 121, 172, 125, 170, 185, 183] enabling approaches that leverage the combination of all these data types. Cross-modal classification is a task that leverages knowledge acquired on one data domain to classify data from another one. This allows, for example, to exploit data domains that are more easily accessible, such as images and audio domains, and apply knowledge on an unexplored domain, like haptic. Cross-modal classification has been performed with audio data in [155], and using visual data in [116, 176, 155, 129]. Furthermore, a subset task of the cross-modal classification application is cross-modal retrieval [156, 157, 121, 125, 180]. Cross-modal retrieval sits between classification and generation applications. The core idea consists of retrieving the data sample in one domain based on a query from another domain by finding the best correspondence within the dataset used to train the system.
To quantitatively evaluate the results obtained and compare them with other works, various metrics have been employed. Other than the standard classification metrics, such as Classification Accuracy, Classification Precision, Recall, F1-score, mean Average Precision (mAP) and rank@, an haptic-specific metric L2 distance aims at measuring the distance between the feature vectors extracted by two instances from different data sources.
@@
refers to the creation of content in one sensory modality based on the information gathered from another one. This process involves translating the haptic information of a surface across different types of media. Recently, this task has gained popularity in many real-world applications, since it allows to collect data in one modality and to transfer the information to different ones. From the haptic perspective, cross-modal generation can be applied to generate vibro-tactile acceleration signals from other types of data, such as texture images [159, 158, 179, 144, 162, 160]. Another approach involves mapping acoustic and acceleration signals, exploiting their similarities [125, 161]. Furthermore, visual and acoustic domains can also be combined [173]. Another data type widely employed in literature for the cross-modal generation task is the haptic map, which enables the usage of computer vision algorithms to map RGB images into haptic data [177, 125, 181, 132, 162, 163]. The most common metrics include Structural Similarity Index Measure (SSIM), PSNR, RMSE, MAE, Fréchet inception distance (FID), and L2-distance, which assess the quality of the generated samples and the difference from the ground truth. In addition to these, many authors employed non-standard metrics. In [181], the authors introduce the Contrastive Visuo-Tactile Pre-Training (CVTP) metric, a cosine similarity that evaluates the distance between visual and tactile embeddings, obtained within a contrastive learning framework. Material Classification Consistency (MCC) [181, 163] is another evaluation strategy that is gaining popularity. This method consists of using a material classifier trained on ground truth samples to classify the cross-generated samples and check whether they correspond to the proper class. In [125], the STFT-distance and Envelope Distance metrics are introduced to evaluate the distance between the ground truth and the spectrograms or signals predicted. Furthermore, in [125], the authors assess the quality of the samples generated within the audio-tactile generation task by taking into account also the perception similarity. To do this, they employed the CDPAM metric, defined in [186].
@@
is a recent application that, given a desired haptic map, aims at ”stylizing” the RGB image of an object, modifying its appearance to reflect the tactile characteristics accurately. This application has been applied mainly to the Touch And Go dataset using methods based on contrastive learning [108] or diffusion models [181, 163]. Despite this task emerging from the use of RGB and Haptic Maps data, the HaTT dataset has also been previously employed. In [145], the authors exploited a physical surface as a constraint, modifying the accelerations emitted by the haptic device to render a different surface according to the model collected in HaTT.
To evaluate the results, two metrics mainly employed. For the Touch and Go dataset, the authors used a material classifier pre-trained on ground truth samples. This approach, previously presented in cross-modal generation tasks, checks if images modified to reflect a specific haptic style are classified correctly as belonging to that style. For the HaTT dataset, the authors verify whether the acceleration generated by the haptic device matched the ground truth model.
@@
is a robotic application that focuses on determining the best way for a robotic hand to manipulate an object, performing tasks like moving or positioning it. While not being strictly modeling-related, it requires the understanding of the stiffness and friction properties of an object, thus modeling of haptic properties plays a fundamental role in this task. These algorithms rely mainly on vision-based sensors to gather haptic information on the part of the object occluded by robotic fingers that cannot be seen by external cameras. To perform grasp prediction various strategies have been developed. In [112, 125], the authors rely on the difference between the initial and final grasp position to train a DNN to predict the success of the grasp. Furthermore, in [117], the authors introduced a conditioning factor that enables the DNN to adjust the grasping action dynamically if the first attempt fails. Starting from the same idea of training the classifier to predict if the grasp will be successful or not, in [164, 165] the authors designed a self-attention mechanism, while in [108, 166, 167, 178, 163] contrastive learning frameworks have been adopted. All these works employ only tactile maps and RGB images to perform grasp prediction. The work in [123] represents the only exception, where, to understand if grasping will be successful in challenging conditions, such as while shaking the object, the authors also employed mono-dimensional signals collected using force/torque sensors to train the classifier. Furthermore, grasp prediction is of particular interest for the works that aim at bridging the gap between simulation and reality, as presented in [125, 167].
For the evaluation, grasp prediction presents well-established metrics: Accuracy, Precision, Recall, and F-score.
@@
is another key application in robotics. This application aims at identifying the exact point on an object where a sample has been taken. Initially, the primary goal of this task was to determine where an object was touched to teach robots how to grasp object correctly [124, 127]. Recently, however, research has moved towards identifying the location on an RGB image or a 3D model where other types of data, such as acoustic signals [125] or haptic signals [132], were generated.
To assess the performances of a contact localization algorithm, the metrics commonly used are mAP, to verify whether the system precision and recall are well balanced, RMSE, to assess the accuracy of both the position and the rotation measurements by comparing them to the ground truth, and average normalized distance (ND), to measure the distance between the predicted and the ground truth points.
@@
is an application aiming at assessing how human touch feels different textures or surfaces. The goal of this task is to identify models capable of predicting human perception based on physical interactions with surfaces. In [146] and [147], the authors propose an experiment using the HaTT dataset where users has to feel different textures by exploring randomly real surfaces made of different materials. The goal of these works is to identify the similarities that different textures present to cluster them according to the feature extracted. In [174], the authors try to cluster the textures, according to the feature perceived by user, into a feature space. To do this they employ respectively a K-means strategy and a DNN method. In [148], the authors exploit GelSight images to capture textures information and, by modeling the user interaction with them, they create virtual textures that close up with their real counterpart.
This task is evaluated by performing user study evaluation and thus its significance strongly depends on the number of participants.
@@
is a very specific application that relies on video streams from vision-based sensors like GelSight or standard RGB cameras. The goal of this task is to retrieve missing frames within videos collected from a specific data source. In [108], two approaches are explored: single-mode, which relies only on previous frames to generate missing ones, and multimodal, which leverages previous frames coming from different data sources. Differently, in [180], the authors rely on a cross-generation framework, using GelSight-captured video information to generate missing frames in RGB videos and vice-versa.
To evaluate the reconstruction capabilities of these methods, the authors used various metrics, such as L1 and L2 distances, and SSIM. Furthermore, in [108], the authors employed also the Learned Perceptual Image Patch Similarity (LPIPS) metric. LPIPS, firstly introduced in [187], aims at assessing the performance of a model by considering also the human perception.
This chapter has explored the different application domains in which haptic datasets have been adopted. The tasks encompassed deal with a multitude of research areas, demonstrating how the research in rapidly evolving in a multidimensional and multimodal fashion, requiring competences in signal processing and representation, machine learning, human perception, to name a few.

### v2

@@ 6 Open Research Directions
Dataset Single mode Multi- modal Cross- modal Cross-modal generation Stylization Compression Grasp prediction Contact localization Perceptual similarity Video generation/ prediction
HaTT [nie2021influence] [liu2023surface] [cai2022multi] [culbertson2017ungrounded] [nie2023compression] [vardar2019fingertip, richardson2022learning] [heravi2024development]
LMT [strese2014haptic] [strese2016multimodal][wei2022alignment] [zheng2016deep, zhang2021partial, zhang2021generative] [zhou2022multimodal, ji2015preprocessing] [liu2022texture, zheng2019cross] [zheng2020lifelong] [li2021research, cai2021visual] [xi2024cm, zhan2023method] [fang2024bidirectional] [hassen2020pvc, zhu2022perceptual] [liu2023online]
FoS [calandra2017feeling][yang2024binding] [cui2020self, liu2023deep] [liu2023self, zhang2023grasp]
Fabrics [wang2019fabric, cao2023learn] [wei2022alignment, zhang2021generative]
Texplorer [strese2019haptic] [strese2017content, strese2019haptic] [khojasteh2023multimodal, liu2018surface] [liu2023surface, wei2021multimodal] [zheng2020lifelong] [song2023cross] [priyadarshini2019perceptnet]
ViTac [luo2018vitac] [cao2020spatio, rouhafzay2020transfer] [rouhafzay2020transfer] [luo2018vitac][rouhafzay2020transfer] [lee2019touching]
MTaF [calandra2018more][dave2024multimodal]
Haptex [cai2022gan] [liu2023online]
VisGel [li2023learning] [yang2023generating] [li2023learning]
ObjectFolder [gao2023objectfolder][li2023vito] [gao2021objectfolder][gao2023objectfolder] [li2023learning, yang2024binding] [gao2023objectfolder] [gao2023objectfolder][yang2024binding] [gao2023objectfolder] [li2023learning]
Touch and Go [dave2024multimodal, lei2024vit] [yang2024binding] [yang2022touch][dave2024multimodal] [lei2024vit, lyu2024omnibind] [yang2024binding, li2023learning] [yang2023generating, yang2024binding] [fang2024bidirectional] [yang2022touch] [yang2024binding, yang2023generating] [yang2022touch][yang2024binding] [yang2022touch][li2023learning]
PoseIt [kanitkar2022poseit]
SSVTP [yang2024binding] [kerr2022self] [yang2024binding] [yang2024binding] [kerr2022self]
MPI-10 [khojasteh2024robust]
SSVTP + HCT [fu2024touch] [fu2024touch]
TaRF [dou2024tactile] [dou2024tactile] [dou2024tactile]
@@
6.1 Haptic surface modeling
In haptic surface modeling, available datasets and modeling methods leave two main research questions open.
What is the best way to collect highly precise haptic data? As shown in Chapter 4, the literature remains divided on the optimal representation and collection of haptic data. On one hand, vision-based data collection—such as images and videos—has gained significant traction in recent years due to its ability to capture spatially distributed surface data and the wide availability of advanced processing techniques in this domain. However, a precise, high-resolution mapping between RGB images and specific spatial points remains elusive, leading to difficulties in estimating tactile features like friction and stiffness. On the other hand, parametric signals can capture highly precise data but struggle to generalize knowledge to neighboring points on a surface. Recent approaches [shin2018geometry, shin2020hybrid, heravi2020learning] have attempted to combine the strengths of both methods, but their applicability is limited by the lack of available datasets. As discussed in Chapter 5, the current literature lacks comprehensive datasets that provide matched haptic maps and one-dimensional signals, which constrains further research developments.
How to validate the modeling quantitatively? Researchers have primarily validated their models through user studies, which, while effective, are difficult and costly to replicate. Another approach involves using the model for downstream tasks like material classification, but this method validates the collected features rather than the model itself. These limitations arise from the inherent challenge of ensuring that the same physical objects are available to all practitioners. One potential solution is to rely on digitally stored 3D models [gao2021objectfolder]. However, these models are often designed either as implicit representations, which only allow the collection of specific signals, or rely on 3D engines that lack tools for incorporating tactile samples into scenes. The availability of a standardized set of shared 3D models, complete with ground-truth data, would enable researchers to quantitatively evaluate and compare different modeling techniques. This would foster the development of a unified validation framework and comparison metrics, ultimately advancing the field.
6.2 Haptic surface processing
In this chapter we provide an entry point for practitioners on the main tasks currently tackled in the haptic signal processing domain and highlight current research directions. Starting from the datasets listed in Section 5, in Tab. 4 we provide a thorough classification of the available works for each task. While some of them, like classification and compression, are akin to other domains and straightforward to understand, other tasks are very specific to the haptic domain, like contact localization and grasp prediction. As shown in the previous paragraph, processing techniques are deeply intertwined with modeling, as they are used as downstream tasks to validate the data modeling, or, vice-versa, they provide strict requirements about the information that should be contained in the data.
For a comprehensive overview, in the next paragraphs we delve into each task, highlighting their peculiarities.
@@ 6.2 Haptic surface processing
It is one of the tasks mainly performed when dealing with haptic data. As the collected datasets usually provide ground truths in different domains (e.g. haptic maps, RGB images, audio produced by the object upon contact) for the same sample, classification can be performed according to different strategies: single mode, multi-modal, and cross-modal. Works adopting the single mode strategy aim to classify materials/objects by training and testing in the same domain. For approaches employing only haptic data, this task is usually performed on tactile maps [luo2018vitac, strese2019haptic, wang2019fabric, cao2020spatio, rouhafzay2020transfer, gao2021objectfolder, wei2021multimodal, cao2023learn, gao2023objectfolder, dave2024multimodal, lei2024vit, yang2024binding], since dealing with images enables the usage of established approaches, such as deep neural networks (DNNs). Differently, other works, such as [strese2014haptic, nie2021influence] proposed a material classification method that exploits acceleration data, achieving state-of-the-art results and showing that acceleration data establish a unique fingerprint for each material. Multi-modal approaches employ multiple data sources to classify materials and objects. These works rely mainly on visual and haptic data [zheng2016deep, rouhafzay2020transfer, zhang2021generative, zhang2021partial, kerr2022self, wei2022alignment, yang2022touch, zhou2022multimodal, li2023vito, liu2023surface, dave2024multimodal, dou2024tactile, fu2024touch, lyu2024omnibind, yang2024binding], since their combination allows for the exploitation of both mono-dimensional and multi-dimensional data. However, also audio cues are widely used jointly with haptic data [ji2015preprocessing, strese2016multimodal, strese2017content, liu2018surface, strese2019haptic, gao2021objectfolder, wei2021multimodal, gao2023objectfolder, khojasteh2023multimodal, khojasteh2024robust, lei2024vit] enabling approaches that leverage the combination of all these data types. Cross-modal classification is a task that leverages knowledge acquired on one data domain to classify data from another one. This allows, for example, to exploit data domains that are more easily accessible, such as images and audio domains, and apply knowledge on an unexplored domain, like haptic. Cross-modal classification has been performed with audio data in [liu2022texture], and using visual data in [luo2018vitac, rouhafzay2020transfer, liu2022texture, fu2024touch]. Furthermore, a subset task of the cross-modal classification application is cross-modal retrieval [zheng2019cross, zheng2020lifelong, gao2021objectfolder, gao2023objectfolder, li2023learning]. Cross-modal retrieval sits between classification and generation applications. The core idea consists of retrieving the data sample in one domain based on a query from another domain by finding the best correspondence within the dataset used to train the system.
@@
It refers to the creation of content in one sensory modality based on the information gathered from another one. This process involves translating the haptic information of a surface across different types of media. Recently, this task has gained popularity in many real-world applications, since it allows to collect data in one modality and to transfer the information to different ones. From the haptic perspective, cross-modal generation can be applied to generate vibro-tactile acceleration signals from other types of data, such as texture images [cai2021visual, li2021research, cai2022gan, cai2022multi, fang2024bidirectional, xi2024cm]. Another approach involves mapping acoustic and acceleration signals, exploiting their similarities [gao2023objectfolder, zhan2023method]. Furthermore, visual and acoustic domains can also be combined [song2023cross]. Another data type widely employed in literature for the cross-modal generation task is the haptic map, which enables the usage of computer vision algorithms to map RGB images into haptic data [lee2019touching, gao2023objectfolder, yang2023generating, dou2024tactile, fang2024bidirectional, yang2024binding].
@@
It is a recent application that, given a desired haptic map, aims at ”stylizing” the RGB image of an object, modifying its appearance to reflect the tactile characteristics accurately. This application has been applied mainly to the Touch And Go dataset using methods based on contrastive learning [yang2022touch] or diffusion models [yang2023generating, yang2024binding]. A method exploiting height maps instead of haptic maps is presented in [tymms2020appearance]. Here the authors define an optimization process that aims at changing the roughness of an object while preserving its visual appearance. Despite this task emerging from the use of RGB and Haptic Maps data, the HaTT dataset has also been previously employed. In [culbertson2017ungrounded], the authors exploited a physical surface as a constraint, modifying the accelerations emitted by the haptic device to render a different surface according to the model collected in HaTT. A different approach employing a perceptual metric is presented in [piovarvci2020fabrication]. Here, the authors use a data-driven method to properly tune the haptic feedback given by digital drawing tools (i.e., drawing surface and stylus), defining different drawing styles.
Compression
It is a key task in the field of signal processing, with the goal of enabling efficient transmission of haptic data through compression and reconstruction strategies. This is particularly important in scenarios requiring real-time communication, such as in medical applications. Research in this area has primarily focused on vibro-tactile signals and has explored both traditional signal processing techniques and deep neural network (DNN)-based approaches. In [hassen2020pvc] they proposed a vibro-tactile perceptual coding method that utilizes a sparse linear prediction strategy. To validate the reconstruction efficiency objectively, the authors exploited the ST-SIM metric, previously introduced in [hassen2019vibrotactile, hassen2019subjective]. This approach has been also integrated into the tactile codec standard by the IEEE1918.1.1 Haptic Codecs Task Group for the Tactile Internet. In another work [nie2023compression], they convert haptic signals into grayscale images, which are then compressed using conventional still image compression techniques. On the DNN side, in [zhu2022perceptual] they introduced a transformer-based architecture for perceptual compression of haptic data, while in [liu2023online] they combined stacked autoencoders with gated recurrent units for online compression.
@@
This is a key task in robotics, to instruct a robotic hand to manipulate an object, performing tasks like moving or positioning it. While not being strictly modeling-related, it requires the understanding of the stiffness and friction properties of an object, thus modeling of haptic properties plays a fundamental role in this task. These algorithms rely mainly on vision-based sensors to gather haptic information on the part of the object occluded by robotic fingers that cannot be seen by external cameras. To perform grasp prediction various strategies have been developed. In [calandra2017feeling, gao2023objectfolder], the authors rely on the difference between the initial and final grasp position to train a DNN to predict the success of the grasp. Furthermore, in [calandra2018more], the authors introduced a conditioning factor that enables the DNN to adjust the grasping action dynamically if the first attempt fails. Starting from the same idea of training the classifier to predict if the grasp will be successful or not, in [cui2020self, liu2023deep] the authors designed a self-attention mechanism, while in [yang2022touch, liu2023self, zhang2023grasp, dave2024multimodal, yang2024binding] contrastive learning frameworks have been adopted. All these works employ only tactile maps and RGB images to perform grasp prediction. The work in [kanitkar2022poseit] represents the only exception, where, to understand if grasping will be successful in challenging conditions, such as while shaking the object, the authors also employed mono-dimensional signals collected using force/torque sensors to train the classifier. A different approach is taken by the method of [lau2016tactile], in which the authors employ a prior knowledge of the mesh of an object to identify the salient points at which a successful grasp can be made. Furthermore, grasp prediction is of particular interest for the works that aim at bridging the gap between simulation and reality, as presented in [gao2023objectfolder, zhang2023grasp]. An alternative approach is proposed in [sundaram2019learning], where the authors developed a dataset of user grasps recorded using a sensor-equipped glove. A neural network was then used for object identification, weight prediction, hand pose estimation, and analyzing hand region correlations during grasping.
@@
It is another key application in robotics. This application aims at identifying the exact point on an object where a sample has been taken. Initially, the primary goal of this task was to determine where an object was touched to teach robots how to grasp object correctly [kerr2022self, suresh2023midastouch]. Recently, however, research has moved towards identifying the location on an RGB image or a 3D model where other types of data, such as acoustic signals [gao2023objectfolder] or haptic signals [dou2024tactile], were generated.
@@
It is an application aiming at assessing how human touch feels different textures or surfaces. The goal of this task is to identify models capable of predicting human perception based on physical interactions with surfaces. In [vardar2019fingertip] and [richardson2022learning], the authors propose an experiment using the HaTT dataset where users has to feel different textures by exploring randomly real surfaces made of different materials. The goal of these works is to identify the similarities that different textures present to cluster them according to the feature extracted. In [priyadarshini2019perceptnet], the authors try to cluster the textures, according to the feature perceived by user, into a feature space. To do this they employ respectively a K-means strategy and a DNN method. In [heravi2024development], the authors exploit GelSight images to capture textures information and, by modeling the user interaction with them, they create virtual textures that close up with their real counterpart.
@@
It is a very specific application that relies on video streams from vision-based sensors like GelSight or standard RGB cameras. The goal of this task is to retrieve missing frames within videos collected from a specific data source. In [yang2022touch], two approaches are explored: single-mode, which relies only on previous frames to generate missing ones, and multimodal, which leverages previous frames coming from different data sources. Differently, in [li2023learning], the authors rely on a cross-generation framework, using GelSight-captured video information to generate missing frames in RGB videos and vice-versa.
This chapter has explored the diverse application domains where haptic datasets have been utilized. The tasks analyzed highlight the need for an interconnected and multimodal research approach to address open challenges in the field. Building on this foundation, major research efforts should focus on improving the realism and adaptability of haptic models to better capture fine-grained textures and dynamic properties of real-world surfaces. Current methods are typically based on predefined sets of materials, but future work should prioritize adaptive frameworks that can interpret scenario information in real time so that they can generalize for any type of surface. This shift introduces not only a data collection problem, but also a scalability challenge, as many existing architectures struggle to run efficiently in runtime environments, posing limitations for robotic and extended reality applications. Finally, as haptic research advances, establishing reliable, objective evaluation metrics becomes crucial. While defining such metrics may be more straightforward in robot-centric applications, scenarios involving human-in-the-loop present additional complexities. The inherent bias in subjective evaluations can undermine standardization efforts, making it difficult to develop universally accepted assessment frameworks.
6.3 Interoperability, implementation and integration
Beyond specific technical challenges, haptics research is increasingly influenced by a number of issues that cut across domains and applications. This section outlines the main limitations in standardisation, multimodal integration, real-time capabilities and evaluation practices that need to be addressed to advance the field.
Missing Standardization. As discussed in Section 3, one of the major limitations hindering cohesive and scalable research in haptics is the absence of a standardized processing pipeline. This issue is twofold: on the one hand, there is no universally accepted data type or format that has proven clearly superior in encoding haptic information. On the other hand, existing datasets lack the scale (in terms of sample size) and diversity (in terms of class variety) that have led to breakthroughs in other domains, such as ImageNet in computer vision. This not only limits the range of features and representations that can be explored and benchmarked, but also limits the effectiveness of data-driven and machine learning frameworks.
Multimodality. The integration of multiple data modalities has been a driving force in areas like vision and language. In haptics, multimodality can be both intra-domain and cross-domain. Intra-domain multimodality currently lacks a unified framework to relate different haptic data types—such as haptic maps and mono-dimensional signals—and to bridge human- and robot-based sensing. Enabling seamless conversion between data formats would allow for dataset expansion and deeper convergence across representations. Cross-domain multimodality remains an underexplored avenue: integrating haptic signals with visual and textual information (e.g., in large language or vision-language models) could dramatically increase the availability of training data and enable richer semantic understanding.
Real-Time Performance. At present, only mono-dimensional haptic signals are effectively used in real-time applications such as teleoperation, due to their relatively low data complexity. More complex data representations—such as haptic images or 3D tactile reconstructions—are still far from being captured, processed, and rendered in real time. Future research should focus on improving computational efficiency and designing hardware capable of low-latency acquisition and feedback.
Evaluation and User Experience. There is a pressing need for better evaluation methodologies in haptic research. New metrics should be developed to assess the fidelity, expressiveness, and perceptual quality of haptic modeling algorithms. While physical modeling can use metrics like MSE or L1 distance, these fail in evaluating results based on perception, which is subjective for definition. Promising efforts, such as the metric proposed in [hassen2019vibrotactile, hassen2019subjective], combine objective computation with subjective validation, but this approach is still underdeveloped. To advance the field, new perceptually grounded metrics and standardized testing protocols across devices and contexts are essential for ensuring reproducibility and improving user experience, especially in virtual and remote interactions applications.
6.4 Applications
Future research in the field of haptics will also encompass a wide range of application-driven studies, where the final use cases are well-defined. In this section, we provide a brief summary of such applications, specifically focusing on the fields of extended reality (XR) and medicine. Comprehensive reviews of current applications of haptic technologies can be found in [giri2021application, zhu2022soft, pacheco2024haptic].
Extended Reality. A significant research gap in haptic technologies is the integration of haptic data representations into graphic 3D engine pipelines, as highlighted in studies such as [wakita2008texturebased, kamuro2012haptic, kim2011haptic, costes2018haptic]. Current technologies often rely on custom libraries that are challenging to access and manipulate, with no shared open-source library available—a resource that would significantly benefit researchers. This issue is likely tied to the aforementioned lack of a standardized pipeline and it represents a fundamental open research direction.
Training. Haptic technology is increasingly employed to train expert workers who must operate in hazardous conditions that are challenging to replicate safely. It allows them to practice procedures and techniques in a simulated environment, improving their skills and reducing the risk of errors during real-life procedures.
Telemedicine. Telemedicine leverages information and telecommunication technology to deliver medical services remotely, overcoming geographical barriers between physicians and patients. While telemedicine has largely focused on communication-related aspects [alenoghena2023telemedicine], integrating haptics enhances remote medical examinations by transmitting touch sensations like pressure and vibration. One key aspect to enhance this practice consists in improving the quality of the signal modeled and transmitted. For example, a haptic device attached to a surgeon’s gloves can simulate tactile feedback during a robot-assisted surgery, allowing the surgeon to feel tissue resistance and pressure, thus improving the quality of care.
Rehabilitation. Haptic technology is employed in applications like prosthetics to provide a more natural and intuitive sense of touch. This enables amputees to achieve greater control and enhanced dexterity, leading to increased independence and an improved quality of life. In this context, developing novel and efficient methods for modeling haptic data to enable real-time haptic feedback is crucial for supporting more effective rehabilitation.
Surgery. In minimally invasive and robotic surgeries, haptic technology provides surgeons with an intuitive sense of touch, enabling them to perform procedures with greater precision and control. This can lead to improved surgical outcomes, shorter recovery times, and reduced risk of complications.
Communication. Although an initial protocol for transmitting haptic data has been developed [steinbach2018haptic], optimizing the compression and reconstruction of haptic data remains a critical challenge for various applications (e.g. telemedicine). In this context, as highlighted in [emami2024survey] research remains open and require further exploration.